# Optimizing Retinal Prosthetic Stimuli with Conditional Invertible Neural Networks

### v1

@@ Abstract
Implantable retinal prostheses offer a promising solution to restore partial vision by circumventing damaged photoreceptor cells in the retina and directly stimulating the remaining functional retinal cells. However, the information transmission between the camera and retinal cells is often limited by the low resolution of the electrode array and the lack of specificity for different ganglion cell types, resulting in suboptimal stimulations. In this work, we propose to utilize normalizing flow-based conditional invertible neural networks to optimize retinal implant stimulation in an unsupervised manner. The invertibility of these networks allows us to use them as a surrogate for the computational model of the visual system, while also encoding input camera signals into optimized electrical stimuli on the electrode array. Compared to other methods, such as trivial downsampling, linear models, and feed-forward convolutional neural networks, the flow-based invertible neural network and its conditional extension yield better visual reconstruction qualities w.r.t. various metrics using a physiologically validated simulation tool.
Keywords:
Invertible neural network Retinal prosthesis Neural engineering.
@@ 1 Introduction
Retinal implants aim to restore partial vision to individuals with certain types of visual impairments, particularly those caused by diseases like retinitis pigmentosa (RP) or age-related macular degeneration (AMD). The visual signals captured by an external camera are transmitted to the electrical signals on the implanted electrode array, which bypass damaged photoreceptor cells in the retina and directly stimulate the remaining functional retinal cells. Various approaches exist to implant the prosthesis to different locations, such as epiretinal (facing the ganglion cells), subretinal (behind the photoreceptors), and suprachoroidal (on top of the choroid). Since the simulation model we utilize from the stimulus to the percept is based on the epiretinal implant Argus I/II devices, we limit our findings for epiretinal implants [31], such as Argus I/II [20], EPIRET3 [21, 25], and VLARS [19, 33]. However, the limited resolution and lack of specificity for different ganglion cell types (such as ON and OFF cells) necessitate an advanced algorithm for stimulation optimization [6].
The retinal prosthetic stimulation optimization can be considered as an inverse problem, where the optimized stimulus is achieved with the inverse transformation of the biologically fixed visual pathway. With the help of powerful machine learning technologies, the optimization tasks can be solved, e.g., using an evolutionary heuristic algorithm [26], a reconstruction decoder [28] or reinforcement learning [17]. Based on a physiologically validated simulation tool, pulse2percept [4], a line of deep learning-based approaches is proposed to learn an inverse model [23, 35], to learn a hybrid autoencoder [12], or to learn an optimization encoder with a surrogate visual pathway simulation model in an end-to-end manner [34]. Moreover, preference-based human-in-the-loop Bayesian optimization algorithms are proposed [10, 11], which facilitate the difficulty in determining the patient-individual parameters of an in silico computational model.
Figure 1: (a) A schematic overview of the stimulation optimization for retinal prostheses. (b) A retinal implant with a $9\times 9$ electrode grid on the axon map [5].
@@
As one of the notable early papers of Invertible Neural Networks (INNs), Dinh et al. [9] propose to use INNs for density estimation tasks. Throughout a series of non-volume transformation processes, the normalizing flow [24] (total probability mass) is preserved, which ensures that the resulting distribution remains a valid probability distribution. Conceptually, INNs are naturally suitable to construct an autoencoder [16], as the forward and backward direction of the INN should always form a perfect encoder-decoder pair. Nguyen et al. [22] train an INN autoencoder with an artificial bottleneck, created by setting several elements to zero before the inverse pass. Sorrenson et al. [27] train autoencoders with likelihood on novel flow architectures that do not rely on coupling and are not dimension preserving. Furthermore, Ardizzone et al. first propose an INN variant to solve inverse problems [3] and an extension to INNs with explicit controls, namely Conditional Invertible Neural Networks (cINNs) [2].
@@
In this work, we follow an INN [3] and a cINN approach [2] to learn a computational model for the visual pathway (Fig. 1a) and automatically obtain the inverse mapping from percept to stimulus thanks to the intrinsic feature of invertibility. Furthermore, INNs are generative and can be leveraged to become bijective to solve inverse problems [3]. In addition, by imposing conditioning to the INN, the generated inverse input (optimized stimulus, see Fig. 2) shall follow the desired constraint. We compare the reconstruction quality of the predicted percepts using cINNs to other approaches, including trivial downsampling, a linear model, feed-forward and invertible neural networks, and address the promising optimization for retinal prosthetic stimuli.
@@ 2.1 Stimulation Optimization
We begin by formulating the in silico optimization problem for retinal prosthetic stimuli. Let ${f}$ represent the function of the visual pathway (Fig. 1), which uses the electrical stimulus $\mathbf{x}_{s}$ from the implanted electrode array as input and produces the percept $\mathbf{y}_{p}$ as output. We assume that $f_{\phi}$ can simulate ${f}$, where $\phi$ denotes the parameters of the computational model, such as the Axon Map Model [5]. The desired encoding function $g_{\theta}:\tilde{\mathbf{x}}\mapsto\mathbf{x}_{s}$ from the visual signal (target) $\tilde{\mathbf{x}}$ to the stimulus ${\mathbf{x}_{s}}$ can be found with the following optimization problem:
$\min_{\theta}\mathcal{L}_{\mathrm{MSE}}\left(\tilde{\mathbf{x}},f_{\phi}\left(g_{\theta}\left(\tilde{\mathbf{x}}\right)\right)\right),$ (1)
where the Mean Squared Error (MSE) serves as the objective function $\mathcal{L}_{\mathrm{MSE}}$. We employ a neural network to construct $g_{\theta}$ with learnable parameters $\theta$.
2.2 Invertible Neural Network
Given a random variable $\mathbf{z}=h(\mathbf{x})$ with a known and tractable density, e.g., a spherical multivariate normal distribution $\mathbf{z}\sim\pi(\mathbf{z})=\mathcal{N}(\mathbf{z};\mathbf{0},\mathbf{I}))$, a data point $\mathbf{x}$ from an i.i.d. dataset $\mathcal{D}$ can be generated by $\mathbf{x}=h^{-1}(\mathbf{z})$ for an invertible function $h$. In this work, we use Invertible Neural Networks (INNs) to express $h$. Following the change of variables theorem, the probability density function of such a model, given a data point $\mathbf{x}$, can be written as:
$p(\mathbf{x})=\pi(\mathbf{z}=h(\mathbf{x}))\left|\det\frac{\partial h(\mathbf{x})}{\partial\mathbf{x}^{T}}\right|,$ (2)
where $\mathbf{J}_{h}(\mathbf{x})=\partial h(\mathbf{x})/\partial\mathbf{x}^{T}$ is the Jacobian matrix of $h$ at $\mathbf{x}$.
If a process $s:\mathbf{x}_{D}\in\mathbb{R}^{D}\mapsto\mathbf{y}_{d}\in\mathbb{R}^{d}$ is not bijective with $D>d$, we can introduce a latent random variable $\mathbf{z}_{K}\sim\pi(\mathbf{z}_{K})=\mathcal{N}(\mathbf{z}_{K};\mathbf{0},\mathbf{I}_{K}))$ with $K=D-d$, as proposed by Ardizzone et al. [3]. We then construct an invertible process $h:\mathbf{x}_{D}\mapsto[\mathbf{y}_{d},\mathbf{z}_{K}]$ with $h(\mathbf{x}_{D})=[h_{\mathbf{y}_{d}}(\mathbf{x}_{D}),h_{\mathbf{z}_{K}}(\mathbf{x}_{D})]$ and $h_{\mathbf{y}_{d}}(\mathbf{x}_{D})\approx s(\mathbf{x}_{D})$. Similar to [3], we apply a supervised Mean Squared Error (MSE) loss on $\mathbf{y}_{d}$ with $\mathcal{L}_{\mathrm{MSE}}=\mathbb{E}[(\mathbf{y}_{d}-h_{\mathbf{y}_{d}}(\mathbf{x}_{D}))^{2}]$ and an unsupervised Maximum Mean Discrepancy (MMD) loss [13] in terms of $\mathbf{x}_{D}$ and $\mathbf{z}_{K}$, respectively. The MMD loss enforces two distributions $P_{M}$ and $P_{M^{\prime}}$ to be identical and is formulated as:
$\mathcal{L}_{\mathrm{MMD}}(p_{M},p_{M^{\prime}})=\left(\mathbb{E}_{i,j}[k(\mu_{i},\mu_{j})-2k(\mu_{i},\mu_{j}^{\prime})+k(\mu_{i}^{\prime},\mu_{j}^{\prime})]\right)^{\frac{1}{2}},$ (3)
@@ 2.2 Invertible Neural Network
As suggested in [3, 30], we utilize the inverse multiquadratic kernel $k(\mu,\mu^{\prime})=\left(1+\left\|(\mu-\mu^{\prime})/\sigma\right\|_{2}^{2}\right)^{-1}$. The MMD loss on $\mathbf{x}_{D}$ forces it to match the prior with $\mathcal{L}_{\mathrm{MMD}}(p(\mathbf{x}_{D}),p(h^{-1}([\mathbf{y}_{d},\mathbf{z}_{K}])))$ and is optional. The MMD loss on $\mathbf{z}_{K}$ forces it to match the normal distribution with $\mathcal{L}_{\mathrm{MMD}}(p(h(\mathbf{x}_{D})),p(\mathbf{y}_{d})p(\mathbf{z}_{K}))$. Moreover, we follow [3] to pad $\mathbf{x}_{D}$ to match the dimensionality if necessary. Under the constraint of the invertible network architecture, both $h$ and $h^{-1}$ can be explicitly trained in both directions with a guaranteed invertible relation. Numerous invertible neural network architecture designs have been proposed to ensure that the inverse and the Jacobian determinant are easily computed, such as additive coupling layer [8], affine coupling layer [9], GLOW (including ActNorm and invertible $1\times 1$ convolution) [15]. The implementation details of the architecture in this work are presented in Section 3.3 and visualized in Fig. 8 in Appendix H.
Figure 2: Training and evaluation pipelines for the stimulation optimization with (a) invertible neural networks and (b) conditional invertible neural networks.

### v2

@@ Abstract
Implantable retinal prostheses offer a promising solution to restore partial vision by circumventing damaged photoreceptor cells in the retina and directly stimulating the remaining functional retinal cells. However, the information transmission between the camera and retinal cells is often limited by the low resolution of the electrode array and the lack of specificity for different ganglion cell types, resulting in suboptimal stimulations. In this work, we propose to utilize normalizing flow-based conditional invertible neural networks to optimize retinal implant stimulation in an unsupervised manner. The inherent invertibility of these networks allows us to use them as a surrogate for the computational model of the visual system, while also encoding input camera signals into optimized electrical stimuli on the electrode array. Compared to other methods, such as trivial downsampling, linear models, and feed-forward convolutional neural networks, the flow-based invertible neural network and its conditional extension yield better visual reconstruction qualities w.r.t. various metrics using a physiologically validated simulation tool.
@@ 1 Introduction
Figure 1: (a) Anatomy of the retina with an electrical epiretinal implant. Modified based on Retinal Implant by Mbuerki under CC BY-SA 3.0 [12]. (b) Electrode activation with the axon map model (top, by Beyeler et al. [8] under CC BY 4.0) and a retinal implant with a $9\times 9$ electrode grid on the axon map (bottom, generated with pulse2percept [7] ).
@@
Retinal implants [60, 65] aim to restore partial vision to individuals with certain types of visual impairments, particularly those caused by diseases like retinitis pigmentosa (RP) or age-related macular degeneration (AMD) [33, 61, 48, 25, 42]. The visual signals captured by an external camera are transmitted to the electrical signals on the implanted electrode array, which bypass damaged photoreceptor cells in the retina and directly stimulate the remaining functional retinal cells (Figure 1(a)). Various approaches exist to implant the prosthesis to different locations, such as epiretinal (facing the ganglion cells), subretinal (behind the photoreceptors), and suprachoroidal (on top of the choroid). Since we utilize the Axon Map Model (Figure 1(b)) as the simulation model from the stimulus to the percept, which is based on the epiretinal implant Argus I/II devices [35, 10], we limit our findings for epiretinal implants [56], such as Argus I/II [35, 10], EPIRET3 [38, 45], and VLARS [59, 34]. However, the limited resolution and lack of specificity for different ganglion cell types (such as ON and OFF cells) necessitate an advanced algorithm for stimulation optimization [47, 23, 28, 36]. The retinal prosthetic stimulation optimization can be considered as an inverse problem, where the optimized stimulus is achieved with the inverse transformation of the biologically fixed visual pathway (Figure 2). With the help of powerful machine learning and deep learning techniques [9, 5], the optimization tasks can be solved, e.g., using an evolutionary heuristic algorithm [46], a reconstruction decoder [55] or reinforcement learning [31]. Based on a physiologically validated simulation tool, pulse2percept [7], a line of deep learning-based approaches are proposed to learn an inverse model [43, 63], to learn a hybrid autoencoder [19], or to learn an optimization encoder with a surrogate visual pathway simulation model in an end-to-end manner [62]. Besides, preference-based human-in-the-loop Bayesian optimization algorithms are proposed [16, 18], which facilitate determining the patient-individual parameters of an in silico retinal prosthetic computational model.
@@
As one of the notable early papers of Invertible Neural Networks (INNs), Dinh et al. [14] propose to use INNs for density estimation tasks. Throughout a series of non-volume transformation processes, the normalizing flow [44] (total probability mass) is preserved, which ensures that the resulting distribution remains a valid probability distribution. Conceptually, INNs are naturally suitable to construct an autoencoder [30], as the forward and backward direction of the INN should form a perfect encoder-decoder pair (but not always [6]). Nguyen et al. [40] train an INN autoencoder with an artificial bottleneck, created by setting several elements to zero before the inverse pass. Sorrenson and Draxler et al. [50] train autoencoders with likelihood on novel flow architectures that do not rely on coupling and are not dimension preserving. Furthermore, Ardizzone et al. first propose an INN variant to solve inverse problems with the maximum mean discrepancy loss (INN-MMD) [3] and an extension to INNs with explicit controls, namely Conditional Invertible Neural Networks (cINNs) [4, 2]. Analogous to other generative models, cINNs can also be applied to unsupervised data synthesis, such as [15]. Zhou et al. [64] apply flow-based invertible generative models to neural encoding and decoding, transforming natural images and corresponding neural spike trains recorded from salamander retinal ganglion cells (RGC) [41] bidirectionally. One key methodological distinction in our approach is that we directly train the stimulus-percept pairs using a physiologically validated simulation tool [7]. In contrast, the neural signals (spike trains) in [64] are trained to reconstruct visual signals without a strong constraint in the decoder in the salamander RGC dataset [41].
Figure 2: A schematic overview of the stimulation optimization for retinal prostheses. The visual signals are converted directly into electrical signals on the electrodes through an optimization process, constrained by a simulation model of the visual pathway.
@@
In this work, we follow an INN [3] and a cINN approach [4, 2] to learn a computational model for the visual pathway (Figure 2) and automatically obtain the inverse mapping from percept to stimulus thanks to the intrinsic feature of invertibility. Furthermore, INNs are generative and can be implemented to become bijective to solve inverse problems [3]. In addition, by imposing conditioning to the INN, the generated inverse input (optimized stimulus, see Figure 3) shall follow the desired constraint. We compare the reconstruction quality of the predicted percepts using cINNs to other approaches, including trivial downsampling, a linear model, feed-forward and invertible neural networks, and address the promising optimization for retinal prosthetic stimuli.
2 Basics of Invertible Neural Network
Given a random variable $\mathbf{z}=h\left(\mathbf{x}\right)$ with a known and tractable density, e.g., a spherical multivariate normal distribution $\mathbf{z}\sim\pi\left(\mathbf{z}\right)=\mathcal{N}\left(\mathbf{z};\mathbf{0},\mathbf{I}\right)$, a data point $\mathbf{x}$ from an i.i.d. dataset $\mathcal{D}$ can be generated by $\mathbf{x}=h^{-1}(\mathbf{z})$ for an invertible function $h$. Following the change of variables theorem, the probability density function of such a model given a data point $\mathbf{x}$ can be written as:
$p(\mathbf{x})=\pi(h(\mathbf{x}))\left|\det\frac{\partial h(\mathbf{x})}{\partial\mathbf{x}^{T}}\right|=\pi(\mathbf{z})\left|\det\mathbf{J}_{h}(\mathbf{x})\right|,$ (1)
where $\mathbf{J}_{h}(\mathbf{x})=\partial h(\mathbf{x})/\partial\mathbf{x}^{T}$ is the Jacobian matrix of $h$ at $\mathbf{x}$ and $\left|\det\mathbf{J}_{h}(\mathbf{x})\right|$ is the absolute value of the Jacobian determinant. Different names describe this approach from different perspectives of the likelihood (Equation 1): Invertible neural network (INN) points out the intrinsic architecture design of $h$ shall be invertible; Normalizing flow (NF) emphasizes that a data distribution of $\mathbf{x}$ is normalized into a simpler form of $\mathbf{z}$ by a series of transformations (flow); Non-volume preserving (NVP, [14]) specifies that the scaling factor $\left|\det\mathbf{J}_{h}(\mathbf{x})\right|$ of volume elements under the flow transformation is not constrained to be $1$.
With $\mathbf{x}$ and $\mathbf{z}$ sharing the identical dimensionality $\mathbf{x},\mathbf{z}\in\mathbb{R}^{n}$ (cf. Section 3.2), the normalizing flow $h$ can be trained by maximizing likelihood of $\mathbf{x}$, which is equivalent to minimizing the Negative Log-Likelihood (NLL) loss:
$\mathcal{L}_{\mathrm{NLL}}$ $=-\log(p(\mathbf{x}))$ (2)
$=-\log\left(\pi(\mathbf{z})\right)-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (3)
$=-\log\left(\mathcal{N}\left(\mathbf{z};\mathbf{0},\mathbf{I}\right)\right)-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (4)
$=-\log\prod_{i=1}^{n}\left(\frac{1}{\sqrt{2\pi}}e^{-\frac{z_{i}^{2}}{2}}\right)-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (5)
$=\frac{n}{2}\log(2\pi)+\frac{1}{2}\|\mathbf{z}\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (6)
$\simeq\frac{1}{2}\|h\left(\mathbf{x}\right)\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|.$ (7)
The first term in Equation 7 can be easily calculated with the output of the forward normalizing flow. The second term is computationally expensive mainly due to the matrix determinant of the Jacobian. A line of INN architecture designs, e.g., additive coupling [13], affine coupling [14] and Glow coupling [29], has been proposed to utilize an LU decomposition style architecture (Equation 13) to achieve the following properties. (1) Invertibility: The design of the coupling layer ensures that the transformation is invertible, which is crucial for computing the exact likelihood of the data under the transformed distribution. (2) Computational efficiency: Both forward and inverse transformations are computationally efficient, which allows the model to scale to high-dimensional data. (3) Easy Jacobian: The Jacobian matrix of the transformation is triangular, making the computation of its determinant straightforward. The coupling layer is structurally similar to the lifting scheme architecture [51] used in discrete wavelet transform. In an affine coupling layer [14], the input is split into two parts with $\mathbf{x}:=[\mathbf{x}_{A},\mathbf{x}_{B}]$, only one part of which is applied with an affine transformation and the other remains unchanged:
$\begin{cases}\mathbf{z}^{\prime}_{A}=\mathbf{x}_{A}\\ \mathbf{z}^{\prime}_{B}=\mathbf{x}_{B}\odot\exp(s(\mathbf{x}_{A}))+t(\mathbf{x}_{A})\,,\end{cases}$ (8)
where $\odot$ denotes the element-wise multiplication, while $s(\cdot)$ and $t(\cdot)$ are scale and translation functions, respectively. $\mathbf{z}^{\prime}:=[\mathbf{z}^{\prime}_{A},\mathbf{z}^{\prime}_{B}]$ represents the intermediate output of the forward one-sided transformation. The invertibility of the first half of the affine coupling layer is straightforward and efficient due to its structure with:
$\begin{cases}\mathbf{x}_{A}=\mathbf{z}^{\prime}_{A}\\ \mathbf{x}_{B}=\left(\mathbf{z}^{\prime}_{B}-t(\mathbf{x}_{A})\right)\odot\exp(-s(\mathbf{x}_{A}))\,.\end{cases}$ (9)
Note that $s(\cdot)$ and $t(\cdot)$ are not required to be invertible themselves and can be any differentiable function (to facilitate gradient-based optimization) including neural networks. The Jacobian matrix of this forward one-sided coupling transformation from $\mathbf{x}$ to $\mathbf{z}^{\prime}$ is triangular, making its log-determinant easy to compute:
$\log\left|\det(\mathbf{J})\right|$ $=\log\left|{\begin{array}[]{cccc}\mathbf{I}&\mathbf{0}\\ \frac{\partial\mathbf{z}^{\prime}_{B}}{\partial{\mathbf{x}_{A}}}&\text{diag}\left({\exp\left({s\left(\mathbf{x}_{A}\right)}\right)}\right)\\ \end{array}}\right|$ (12)
$=\sum s({\mathbf{x}_{A}}).$ (13)
Additionally, in the subsequent layer, the second segment remains unchanged, whereas the first segment will be transformed with the intermediate output from the other side:
$\begin{cases}\mathbf{z}_{A}=\mathbf{z}^{\prime}_{A}\odot\exp(s(\mathbf{z}^{\prime}_{B}))+t(\mathbf{z}^{\prime}_{B})\\ \mathbf{z}_{B}=\mathbf{z}^{\prime}_{B}\,.\end{cases}$ (14)
More details about the INN architectures used in this paper can be found in Section 4.3 and in Figure 4.
The original generative paradigm (e.g., RealNVP [14]) is not suitable for the application of retinal stimulus optimization, as the optimized stimulus shall be guided by the target visual signal as the reference. Therefore, we utilize two INN variants, namely INN-MMD [3] with the combination of an unsupervised MMD loss (Section 3.2) and a supervised MSE loss, and cINN [4, 2] with an explicit conditioning mechanism (Section 3.3).
(a) INN-MMD
(b) cINN
Figure 3: Training and evaluation pipelines for the stimulation optimization with (a) invertible neural networks with maximum mean discrepancy loss (INN-MMD) [3, 4] and (b) conditional invertible neural networks (cINN) [2].
@@ 3.1 Stimulation Optimization
We begin by formulating the in silico optimization problem for retinal prosthetic stimuli. Let ${f}$ represent the function of the visual pathway (Figure 2), which uses the electrical stimulus $\mathbf{x}_{s}$ from the implanted electrode array as input and produces the percept $\mathbf{y}$ as output. We assume that $f_{\phi}$ can simulate ${f}$, where $\phi$ denotes the parameters of the computational model, such as the Axon Map Model [8]. The desired encoding function $g_{\theta}:\tilde{\mathbf{x}}\mapsto\mathbf{x}_{s}$ from the visual signal (target) $\tilde{\mathbf{x}}$ to the stimulus ${\mathbf{x}_{s}}$ can be found with the following optimization problem:
[…]
where $\mathcal{L}_{\mathrm{eval}}$ serves as the objective function. We employ a neural network to construct $g_{\theta}$ with learnable parameters $\theta$. In practice, we often train a differentiable neural network as a surrogate for $f_{\phi}$ (e.g., [62, 19]), enabling convenient end-to-end optimization for the encoder $g_{\theta}$.
Figure 4: We follow [4, 2] to design the INN and cINN architectures with Glow coupling layers [29]. The conditional input $\mathbf{c}$ in both forward and inverse Glow coupling blocks is the only difference between the INN and the cINN approach.
3.2 Invertible Neural Network for Inverse Problems with Maximum Mean Discrepancy
If a process $q:\mathbf{x}_{D}\in\mathbb{R}^{D}\mapsto\mathbf{y}_{d}\in\mathbb{R}^{d}$ is not bijective with $D>d$, we can introduce a latent random variable $\mathbf{z}_{K}\sim\pi\left(\mathbf{z}_{K}\right)=\mathcal{N}\left(\mathbf{z}_{K};\mathbf{0},\mathbf{I}_{K}\right)$ with $K=D-d$, as proposed by Ardizzone et al. [3]. We then construct an invertible process $h:\mathbf{x}_{D}\mapsto[\mathbf{y}_{d},\mathbf{z}_{K}]$ with $h(\mathbf{x}_{D})=[h_{\mathbf{y}_{d}}(\mathbf{x}_{D}),h_{\mathbf{z}_{K}}(\mathbf{x}_{D})]$ and $h_{\mathbf{y}_{d}}(\mathbf{x}_{D})\approx q(\mathbf{x}_{D})$. Similar to [3], we apply a supervised Mean Squared Error (MSE) loss on $\mathbf{y}_{d}$ with $\mathcal{L}_{\mathrm{MSE}}=\mathbb{E}[(\mathbf{y}_{d}-h_{\mathbf{y}_{d}}(\mathbf{x}_{D}))^{2}]$ and an unsupervised Maximum Mean Discrepancy (MMD) loss [22, 21] in terms of $\mathbf{x}_{D}$ and $\mathbf{z}_{K}$, respectively. The MMD loss enforces two distributions $p_{M}$ and $p_{M^{\prime}}$ to be identical and is formulated as:
$\begin{split}\mathcal{L}_{\mathrm{MMD}}(p_{M},p_{M^{\prime}})=&\\ \left(\mathbb{E}_{i,j}[k(\mu_{i},\mu_{j})\right.&\left.-2k(\mu_{i},\mu_{j}^{\prime})+k(\mu_{i}^{\prime},\mu_{j}^{\prime})]\right)^{\frac{1}{2}},\end{split}$ (16)
@@ 3.2 Invertible Neural Network for Inverse Problems with Maximum Mean Discrepancy
As suggested in [3, 53], we utilize the radial basis function-based (RBF) rational quadratic kernel [39] (also known as inverse multiquadric kernel):
$k(\mu,\mu^{\prime})=\left(1+\left\|(\mu-\mu^{\prime})/\sigma\right\|_{2}^{2}\right)^{-1}\,.$ (17)
The MMD loss on $\mathbf{x}_{D}$ forces it to match the prior with $\mathcal{L}_{\mathrm{MMD}}(p(\mathbf{x}_{D}),p(h^{-1}([\mathbf{y}_{d},\mathbf{z}_{K}])))$ and is optional. The MMD loss on $\mathbf{z}_{K}$ forces it to match the normal distribution with $\mathcal{L}_{\mathrm{MMD}}(p(h(\mathbf{x}_{D})),p(\mathbf{y}_{d})p(\mathbf{z}_{K}))$. Moreover, following the approach in [3], we pad $\mathbf{x}_{D}$ with zeros to match the required dimensionality when necessary. This allows us to train both directions as the forward pass, as illustrated in Figure 5. Under the constraint of the invertible network architecture, both $h$ and $h^{-1}$ can be trained in both directions with a guaranteed invertible relation. Numerous invertible neural network architecture designs have been proposed to ensure that the inverse and the Jacobian determinant are easily computed, such as additive coupling layer [13], affine coupling layer [14] and Glow coupling layer [29] (see Section 2). The implementation details of the architecture in this work are visualized in Figure 4.
Table 1: Summary of notations for retinal implants (Section 3.1) in the context of invertible neural networks.
Notations INN-MMD [3] cINN [4, 2]
$f_{\phi}$ Forward network ($h$)
$g_{\theta}$ Reverse network ($h^{-1}$)
$\mathbf{x}_{s}$ Forward input ($\mathbf{x}$ or $\mathbf{x}_{D}$)
$\mathbf{\tilde{x}}$ Reverse input Reverse conditioning
$\mathbf{y}$, $\mathbf{y}_{d}$ Forward output (MSE) Forward conditioning
$\mathbf{z}$, $\mathbf{z}_{K}$ Forward output (MMD) Forward output (NLL)